# Stress-Testing Long-Context Language Models with Lifelong ICL and Task Haystack

### v1

@@ 5.2 Controlled Analysis on Long-Context Utilization
Previously, we find that long-context LMs struggle in the Task Haystack evaluation. In the following section, we investigate the reasons attributing to their failures with various controlled settings.
@@
We investigate the effect of recency bias by comparing the results of Recall and Replay. By replaying ICL demonstrations immediately before testing, model’s average accuracy improve by 1.6% for Mistral-7B and 2.9% for FILM-7B. Replay can be also considered as an oracle that approximates potential mitigating strategies such as prompting the model to recall relevant information (Shi et al., 2023; Anthropic, 2024). However, the improvements only close about half the gap between Baseline and Recall, suggesting that recency bias contributes to but does not fully explain the performance gap.
@@
We examine the effect of irrelevant context, by contrasting Baseline with Random. The results indicate that prepending an irrelevant long text will influence the performance negatively, which corroborates with recent work investigating the robustness of large language models (Levy et al., 2024). Further, Replay can be seen as prepending a long prefix of mostly irrelevant tasks before performing Single-task ICL (Baseline), and thus the gap between Replay and Baseline may be interpreted as caused by prepending irrelevant contexts.
@@
We further compare Baseline, Random, Repeat settings altogether, where Random introduces irrelevant context and Repeat includes only relevant context. Perhaps surprisingly, performance drops in the Repeat setting (-1.3% for Mistral-7B and -3.1% for FILM-7B), where both distractions and recency biases are absent. This observation raises concerns on whether longer inputs are more likely to trigger failure modes and give rise to undesired behaviors in general. While more evidence is needed to derive a conclusion, we suggest that long-context LM users be cautious about including everything in the context window, and recommend using external filtering mechanisms like retrieval augmentation given these undesired behaviors.
@@
We conduct further investigation with the Random, Repeat, Repeat+Shuffle setting, by varying the size of the context and the number of repetitions. Results are reported in Fig. 5. Interestingly, model performance first increases and then dips when running in-context learning for multiple “epochs”. One direct takeaway is that repeating the ICL examples multiple times can potentially improve performance, which may have practical utilities in certain low-data high-inference-budget regimes. However, model performance starts to degrade after repeating more than 8 times. This phenomenon can be interpreted in two ways: (1) It is a known issue that repetition may lead to model degeneration (Nasr et al., 2023); Repeat+Shuffle can possibly alleviate this issue by introducing slight variations in each repeat, which explains that in general Repeat+Shuffle outperforms Repeat. (2) It is also possible that the model “overfits” to the few-shot training data after multiple “epochs”, analogous to the common observations in gradient-based fine-tuning. We invite future work to further investigate the working mechanism of ICL in this “multi-epoch” setting.
@@ 5.3 Additional Observations and Analysis
While examining Task Haystack results, we find that the passing and failing behaviors are highly task-specific. For example, in Fig. 22, Mistral-7B (32k) fails on news_data and insincere_questions in all permutations, meanwhile passes on more popular tasks like boolq and yahoo_answer_topics. We hypothesize that models may have memorized some of the tasks during pre-training or post-training, making these tasks less subjective to performance drop in Lifelong ICL. Alternatively, a task may be too challenging for the model to learn through ICL, and thus it passes the test by maintaining low performance in both Single-task ICL and Lifelong ICL.
To account for these situations, we split all tasks into 2 groups for each model. Tasks of which 4-shot performance is significantly better than 1-shot performance are classified as ICL-effective tasks, and the remaining tasks are considered to be ICL-ineffective. We report the pass rates for each model on these two groups in Table 4. For 10 out of 12 models, pass rates on ICL-effective tasks are lower than pass rate on ICL-ineffective tasks, suggesting that these models tend to “forget” tasks that are newly acquired, and that the overall pass rate may be an overestimate.
@@
Llama2-7B (32k) 6 33.3 10 46.0 41.2 Yi-9B (200k) 6 50.0 10 72.0 63.7
Llama2-7B (80k) 3 80.0 13 100.0 96.3 Yi-34B (200k) 3 46.7 13 67.7 63.8
Llama3-8B (1048k) 6 40.0 10 90.0 71.3 GPT-3.5-Turbo (16k) 5 44.0 11 70.9 62.5
Llama3-70B (1048k) 4 35.0 12 65.0 57.5 GPT-4o (128k) 6 96.7 10 76.0 83.7
@@
While our study mainly focus on undesired performance degradation in Lifelong ICL, which is analogous to catastrophic forgetting in lifelong learning, we also observe trends of positive forward and backward transfers, two desired properties of lifelong learning. [^6] Forward transfer occurs when “a model reuses knowledge acquired from previous tasks when learning new ones”; backward transfer refers the the phenomenon that “a model achieves improved performance on previous tasks after learning a new task.” (Biesialska et al., 2020) In our pass rate design, we deliberately choose two-sided t-test to account for both performance gains and drops. We observe positive transfers in Fig. 3, represented by the blue-colored cells in the 1-shot (4k) column and the last row (94% depth). Similar observations can be made with Llama-2 (32k) in Fig. 15 and GPT-4o in Fig. 20. Additionally, Mistral-7B achieves +3.4% performance gain in the Remove setting compared to the Zero-shot baseline (Fig. 5). We consider these as initial evidence for positive transfer in Lifelong ICL, and invite more rigorous analysis to further explore the properties of Lifelong ICL.
@@ 6 Discussion
(1) As an initial exploration in the Lifelong ICL setting, we primarily focuses on English-only text classification tasks. This potentially limits a comprehensive assessment of model capabilities across various challenges. To get a more complete picture, the evaluation suite may be improved by involving a wider variety of tasks categories (e.g., question answering, conditional generation), modalities (e.g., vision, speech), and languages. We encourage future research to build upon our foundation and explore these more comprehensive settings. (2) This work simplifies the lifelong learning stream by assuming a sequential order, clear task boundaries, and a fixed number of examples per class for each task. Real-world scenarios likely involve a more dynamic learning stream, without clear task boundaries or assumptions on the sequential order. Future work may explore more realistic training streams that reflect this complexity. (3) Finally, due to computational constraints, our evaluation utilizes 5 random permutations of tasks order and 5 different random samples of few-shot training sets. Experimenting with a larger number of samples could potentially reduce the randomness inherent in the results and increase the reliability of the findings. Additionally, we limit our evaluation to up to 32k input tokens. Stress-testing long-context models with their full context lengths may reveal further limitations.
@@
This work leverages openly available datasets that were carefully reviewed by the authors to mitigate potential data privacy and security concerns. We emphasize that these datasets are used solely for evaluation purposes. As our research does not involve direct model weight optimization or the release of new models, the risk of amplifying biases within the data is minimal.
@@ 7 Conclusion
In this paper, we introduced Lifelong ICL, a novel problem setting for long-context language models, and developed Task Haystack, a concrete evaluation suite focusing on evaluating and diagnosing long-context LMs in the Lifelong ICL setting. Our experiments with 12 long-context LMs revealed that while these models excel at needle-in-a-haystack style evaluation, their ability to utilize the context flexibly and contextually remains limited. Through our controlled analysis, we dissected and quantified factors such as recency biases and distractions that contribute to performance drops. We also identified performance degradation when repeating ICL examples or using paraphrased instructions, highlighting a fundamental vulnerability in current long-context models.
Our results demonstrate that Task Haystack poses significant challenges for current top-ranked long-context models. We hope that Lifelong ICL and Task Haystack will serve as valuable tools for diagnosing and advancing the development of future long-context LMs. Additionally, we consider our work as an exploratory step towards backprop-free algorithms in lifelong learning settings.
@@ Acknowledgment
We would like to thank Xisen Jin, Jun Yan, Ting-Yun Chang, Daniel Firebanks-Quevedo for insightful discussions. This work was supported by Cohere For AI Research Grant Program and OpenAI Researcher Access Program. Qinyuan Ye was supported by a USC Annenberg Fellowship.
@@ A.1 Models
We list the details of open-weighted models evaluated in Table 5. For closed models from OpenAI, the specific model versions we evaluated are gpt-3.5-turbo-0125 and gpt-4o-2024-05-13.
Table 5: Open-weight models evaluated in this work.
@@
Llama2-7B 32k TogetherAI [2024] togethercomputer/LLaMA-2-7B-32K
Llama2-7B 80k Fu et al. [2024] yaofu/llama-2-7b-80k
Llama3-8B 1048k GradientAI [2024b] gradientai/Llama-3-8B-Instruct-Gradient-1048k
Llama3-70B 1048k GradientAI [2024a] gradientai/Llama-3-70B-Instruct-Gradient-1048k
@@ A.2 Tasks
We use publicly available datasets in Task Haystack. We provide their references and huggingface identifiers in Table 7. For further use, readers should refer to the licenses of the original datasets.
A.3 Problem Setting Details
Definition of “Pass Rate”.
In §3.2 we introduced “Pass Rate” as the core evaluation metric in Task Haystack. Here we further explain its definition and our considerations when desigining this metric. As illustrated in Fig. 6, we first obtain two groups of 5 different performance metrics, one group using Lifelong ICL prompts, and one group using Single-task ICL prompts; we then use two-sided t-test to examine whether the two groups are significantly different. More specifically, we use scipy.stats.ttest_rel that returns the t-statistic and p-value for the test, and we consider tests with $p<0.05$ as significant differences. Additionally, we choose two-sided test to account for potential positive transfers that may arise in the Lifelong ICL setting (§5.3).
@@ A.3 Problem Setting Details
A.4 Implementation and Engineering Details
Data Preprocessing.
1{
2 "name": "ag_news",
3 "task_type": "classification",
4 "options": ["World", "Sports", "Business", "Technology"],
[…]
9}
[…]
LLM Inference.
We did not use model-specific prompts (e.g., chat template, special tokens, instructions optimized for a specific model). This decision reduces experiment complexity and is reasonable because: (1) we expect a model optimized for chat to still be able to perform ICL as a text-token prediction task; (2) the special tokens may create tokenization inconsistencies in in-context learning; (3) Task Haystack is based on a comparative assumption, making absolute accuracies less important.
Inference Costs.
Running a 64-task, 2-shot Task Haystack experiment with a 7B model on one A6000 GPU takes around 20 hours. Running a 16-task, 8-shot Task Haystack experiment with a 7B model on one A6000 GPU takes around 8 hours. For 34B and 70B models, we use four A6000 GPUs. For evaluations with OpenAI models, we use the Batched API. [^7] https://platform.openai.com/docs/guides/batch All experiments using OpenAI models (Table 2 bottom rows; Fig 20-20) incur a total cost of about \$8,000.
Appendix B Additional Results
B.1 Scale-Task Experiments
In Table  6 we report the results in the Scale-Task setting, where we fix the number of shots per class $n_{shot}$ to be 2, and experiment with $n_{task}\in\{8,16,24,32,40,48,56,64\}$. [^8] Since each column in the table uses a different set of tasks, accuracies and pass rates from different columns are not directly comparable. We noticed that the overall pass rates are higher than those in the Scale-Shot setting (Table 2), potentially due to a smaller value of $n_{shot}$. However, long-context models still struggle in this setting: in 37 out of 40 cases in Table 6, the overall pass rates drop below 90%. The three cases achieving pass rates above 90% use the Llama2-7B (80k) model, an outlier model §5.1.
Table 6: Main Results: Fixing 2 Shots, Scaling the Number of Tasks. See the caption of Table 2 for the explanations of the table headers.
Model 8 tasks (4k) 16 tasks (8k) 32 tasks (15k) 64 tasks (25k)
s-acc l-acc pass s-acc l-acc pass s-acc l-acc pass s-acc l-acc pass
FILM-7B (32k) 79.1 77.1 87.5 79.1 75.1 77.5 73.3 72.0 88.1 70.6 69.7 75.3
Llama2-7B (32k) 70.1 60.7 65.0 72.8 64.5 53.8 70.6 64.5 59.4 67.1 61.2 63.1
Llama2-7B (80k) 49.9 58.5 97.5 49.8 60.2 100.0 49.5 58.3 91.2 48.6 52.0 89.7
Llama3-8B (1048k) 68.3 65.4 75.0 70.0 69.1 76.2 67.4 65.1 75.6 66.4 65.7 81.2
Llama3-70B (1048k) 77.1 73.8 45.0 79.0 74.4 50.0 76.0 61.6 59.4 74.1 70.4 70.3
Yi-6B (200k) 72.0 54.4 50.0 73.0 58.6 51.2 68.4 59.2 63.7 63.7 55.7 65.6
Yi-9B (200k) 78.6 73.4 62.5 77.7 72.9 71.2 75.5 70.3 61.3 70.2 66.8 61.3
Yi-34B (200k) 66.1 70.7 87.5 74.1 72.4 60.0 74.0 69.7 63.1 71.5 68.2 59.4
B.2 Controlled Analysis
In Fig. 8-8 we repeat the controlled experiments in Fig. 5-5, using N-task=64, N-shot=2 instead of N-task=16, N-shot=4. Our observations are generally consistent with those in Fig. 5-5. One exception is Mistral-7B (32k) experiments in Fig. 8, where the model achieves comparable accuracies in Recall, Replay and Remove. We attribute this to the usage of a smaller N-shot value compared to Fig. 5.
@@ B.2 Controlled Analysis
Appendix C NIAH-style Visualizations
In Fig. 11-20 we present detailed Task Haystack results for ten open-weight models. In Fig. 20-20 we present results for the two OpenAI models, GPT-3.5-Turbo and GPT-4o.

### v2

@@ 5.2 Controlled Analysis on Long-Context Utilization
Previously, we find that long-context LMs struggle in the Task Haystack evaluation. In the following section, we investigate the reasons that contribute to their failures with various controlled analyses.
@@
We investigate the effect of recency bias by comparing the results of Recall and Replay. By replaying ICL demonstrations immediately before testing, model’s accuracy improves by 1.6% for Mistral-7B and 2.9% for FILM-7B. Replay can be also considered as an oracle for potential mitigating strategies such as prompting the model to recall relevant information (Shi et al., 2023; Anthropic, 2024). However, the improvements only close about half the gap between Baseline and Recall, suggesting that recency bias contributes to but does not fully explain the performance gap.
@@
We examine the effect of irrelevant context, by contrasting Baseline with Random. The results indicate that prepending an irrelevant long text will influence the performance negatively, which corroborates with recent work investigating the robustness of language models (Levy et al., 2024). Further, Replay can be seen as prepending a long prefix of mostly irrelevant tasks before performing Single-task ICL (Baseline), and thus the gap between Replay and Baseline may be interpreted as caused by prepending irrelevant contexts.
@@
We further compare Baseline, Random, Repeat settings altogether, where Random introduces irrelevant context and Repeat includes only relevant context. Perhaps surprisingly, performance drops in the Repeat setting (-1.3% for Mistral-7B and -3.1% for FILM-7B), where both distractions and recency biases are absent. This observation raises concerns on whether longer inputs are more likely to trigger failure modes and give rise to undesired behaviors in general. While more evidence is needed to derive a conclusion, we suggest that long-context LM users be cautious about including everything in the context window, and we recommend using external filtering or retrieval models when necessary.
@@
We conduct further investigation with the Random, Repeat, Repeat+Shuffle setting, by varying the size of the context and the number of repetitions. Results are reported in Fig. 5. Interestingly, model performance first increases and then dips when running in-context learning for multiple “epochs.” One direct takeaway is that repeating the ICL examples multiple times can potentially improve performance, which may have practical utilities in certain low-data high-inference-budget regimes. However, model performance starts to degrade after repeating more than 8 times. This phenomenon can be interpreted in two ways: (1) It is a known issue that repetition may lead to model degeneration (Nasr et al., 2023); Repeat+Shuffle can possibly alleviate this issue by introducing slight variations in each repeat, which explains why Repeat+Shuffle outperforms Repeat in general. (2) It is also possible that the model “overfits” to the few-shot training data after multiple “epochs”, analogous to the common observations in gradient-based fine-tuning. We invite future work to investigate the working mechanism of ICL in this “multi-epoch” setting.
@@ 5.3 Additional Observations and Analysis
While examining Task Haystack results, we find that the passing and failing behaviors are highly task-specific. For example, in Fig. 24, Mistral-7B (32k) fails on news_data and insincere_questions in all permutations, meanwhile passes on more popular tasks like boolq and yahoo_answer_topics. We hypothesize that models may have memorized some of the tasks during pre-training or post-training, making these tasks less subjective to performance drop in Lifelong ICL. Alternatively, a task may be too challenging for the model to learn through ICL, and thus it passes the test by maintaining low performance in both Single-task ICL and Lifelong ICL settings.
To account for these situations, we split all tasks into 2 groups for each model. Tasks of which 4-shot performance is significantly better than 1-shot performance are classified as ICL-effective tasks, and the remaining tasks are considered to be ICL-ineffective. We report the pass rates for each model on these two groups in Table 4. For 10 out of 12 models, pass rates on ICL-effective tasks are lower than pass rates on ICL-ineffective tasks, suggesting that these models tend to “forget” tasks that are newly acquired, and that the overall pass rates may be an overestimate.
@@
Llama-2-7B (32k) 6 33.3 10 46.0 41.2 Yi-9B (200k) 6 50.0 10 72.0 63.7
Llama-2-7B (80k) 3 80.0 13 100.0 96.3 Yi-34B (200k) 3 46.7 13 67.7 63.8
Llama-3-8B (1048k) 6 40.0 10 90.0 71.3 GPT-3.5-Turbo (16k) 5 44.0 11 70.9 62.5
Llama-3-70B (1048k) 4 35.0 12 65.0 57.5 GPT-4o (128k) 6 96.7 10 76.0 83.7
@@
While our study mainly focus on undesired performance degradation in Lifelong ICL, which is analogous to the catastrophic forgetting phenomenon in lifelong learning, we also observe trends of positive forward and backward transfers, two desired properties of lifelong learning. [^7] Forward transfer occurs when “a model reuses knowledge acquired from previous tasks when learning new ones”; backward transfer refers the the phenomenon that “a model achieves improved performance on previous tasks after learning a new task.” (Biesialska et al., 2020) In our pass rate design, we deliberately choose two-sided t-test to account for both performance gains and drops. We observe positive transfers in Fig. 3, represented by the blue-colored cells in the 1-shot (4k) column and the last row (94% depth). Similar observations can be made with Llama-2 (32k) in Fig. 13 and GPT-4o in Fig. 22. Additionally, Mistral-7B achieves +3.4% performance gain in the Remove setting compared to the Zero-shot baseline (Fig. 5). We consider these as initial evidence for positive transfer in Lifelong ICL, and invite more rigorous analysis to further explore the properties of Lifelong ICL.
@@ 6 Discussion
(1) As an initial exploration in the Lifelong ICL setting, we primarily focuses on English-only text classification tasks. This potentially limits a comprehensive assessment of model capabilities across various challenges. To get a more complete picture, the evaluation suite may be improved by including more diverse tasks categories (e.g., question answering, conditional generation (Ye et al., 2021)), modalities (e.g., vision (Sharma et al., 2024), speech), and languages. We encourage future research to build upon our foundation and explore these more complex settings. (2) This work simplifies the lifelong learning stream by assuming a sequential order, clear task boundaries, and a fixed number of examples per class for each task. Real-world scenarios likely involve a more dynamic learning stream, without clear task boundaries or assumptions on the sequential order. In §B.6, we conduct preliminary experiments by interleaving examples of multiple tasks in the context. Future work may explore more realistic lifelong learning streams with increased complexity. (3) Finally, due to computational constraints, our evaluation utilizes 5 random permutations of tasks order and 5 different random samples of few-shot training sets. Experimenting with a larger number of samples could potentially reduce the randomness inherent in the results and increase the reliability of the findings. Additionally, we limit our evaluation to up to 32k input tokens. Stress-testing long-context models with their full context lengths may reveal further limitations of these models.
@@
This work leverages openly available datasets that were carefully reviewed by the authors to mitigate potential data privacy and security concerns. To the best of our knowledge, the datasets we use do not contain personally identifiable information. Some datasets contain offensive content when the underlying task is offensive content (e.g., hate speech) classification. We emphasize that these datasets are used solely for evaluation purposes. As our research does not involve model training or the release of new models, the risk of amplifying biases within the data is minimal.
@@ 7 Conclusion
In this paper, we introduced Lifelong ICL, a novel problem setting for long-context LMs, and developed Task Haystack, a concrete evaluation suite focusing on evaluating and diagnosing long-context LMs in the Lifelong ICL setting. Our experiments with 14 long-context LMs revealed that while these models excel at needle-in-a-haystack style evaluation, their ability to utilize the context flexibly and contextually remains limited. Through our controlled analysis, we dissected and quantified factors such as recency biases and distractions that contribute to performance drops. We also identified performance degradation when repeating ICL examples or using paraphrased instructions, highlighting a fundamental vulnerability in current long-context models.
Our results demonstrate that Task Haystack still poses significant challenges for newly-released long-context models. We hope that Lifelong ICL and Task Haystack will serve as valuable tools for diagnosing and advancing the development of future long-context LMs. Additionally, we consider our work as an exploratory step towards backprop-free algorithms in lifelong learning settings.
@@ Acknowledgment
We thank the anonymous reviewers for their thoughtful feedback and active engagement throughout the discussion period. In addition, we thank Xisen Jin, Jun Yan, Ting-Yun Chang, Daniel Firebanks-Quevedo, Johnny Wei, Ryan Wang, Wenbo Zhang for insightful discussions. This work was supported in part by Cohere For AI Research Grant Program and OpenAI Researcher Access Program. Qinyuan Ye was supported by a USC Annenberg Fellowship.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See §6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See §6.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See footnote 1.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See §A.4.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] We did not report error bars directly, but we addressed randomness issues by sampling 5 few-shot samples and utilizing a customized pass rate metric. See §3.2.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See §A.4
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See §A.2.
    2. (b)
      Did you mention the license of the assets? [Yes] See §A.2.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See footnote 1.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See §A.2.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See §6.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ A.1 Models
We list the details of open-weighted models evaluated in Table 5. For closed models from OpenAI, the specific model versions we evaluated are gpt-3.5-turbo-0125 and gpt-4o-2024-05-13. For Gemini 1.5 Flash, we evaluated gemini-1.5-flash-001.
Table 5: Open-weight Long-context LMs Evaluated in This Work.
@@
Llama-2-7B 32k TogetherAI [2024] togethercomputer/LLaMA-2-7B-32K
Llama-2-7B 80k Fu et al. [2024] yaofu/llama-2-7b-80k
Llama-3-8B 1048k GradientAI [2024b] gradientai/Llama-3-8B-Instruct-Gradient-1048k
Llama-3-70B 1048k GradientAI [2024a] gradientai/Llama-3-70B-Instruct-Gradient-1048k
Llama-3.1-70B 128k Dubey et al. [2024] meta-llama/Llama-3.1-70B
@@ A.2 Tasks
We select 64 classification tasks from huggingface datasets [Lhoest et al., 2021], following the desiderata listed in §4. We provide their references and huggingface identifiers in Table 6. For further use, readers should refer to the licenses of the original datasets.
A.3 Details on “Pass Rate”
“Pass Rate” Definition.
In §3.2, we introduced “pass rate” as the core evaluation metric in Task Haystack. Here we further explain its definition and our considerations when designing this metric. As illustrated in Fig. 6, we first obtain two groups of 5 different performance metrics (e.g., accuracies), one group using Lifelong ICL prompts, and one group using Single-task ICL prompts; we then use two-sided t-test to examine whether the two groups are significantly different. More specifically, we use scipy.stats.ttest_rel that returns the t-statistic and p-value for the test, and we consider tests with $p<0.05$ as significant differences. We choose to use two-sided tests to account for potential positive transfers that may arise in the Lifelong ICL setting (§5.3).
@@ A.3 Details on “Pass Rate”
“Pass Rate” Limitations.
When computing and aggregating pass rates for multiple tasks, we are effectively performing multiple t-tests in parallel, which might increase the risk of Type I errors. We acknowledge this as a limitation of our approach. Following reviewer feedback, we have tried Bonferroni Correction and Benjamini-Hochberg Correction to account for this. However, these methods lead to new challenges. The first method significantly increases the risk of Type II errors and may lead to overestimated pass rates. The second method may lead to unfair comparison across models. Given these concerns, we decide to maintain the current design of pass rates. While this may affect the quantitative results, the qualitative conclusions in the paper are not expected to change.
@@
A.4 Implementation and Engineering Details
Data Preprocessing.
1{
2 "name": "ag_news",
3 "task_type": "classification",
4 "options": ["World", "Sports", "Business", "Technology"],
[…]
9}
[…]
LLM Inference.
We did not use model-specific prompts (e.g., chat template, special tokens, instructions optimized for a specific model). This decision reduces experiment complexity and is reasonable because (1) we expect a model optimized for chat to still be able to perform ICL as a text-token prediction task; (2) the special tokens (e.g., <|user|>, [INST]) may create tokenization inconsistencies in in-context learning (e.g., World and _World may be two different tokens in the vocabulary); (3) Task Haystack is based on a comparative assumption, making absolute accuracies less important.
Inference Costs.
Running a 64-task, 2-shot Task Haystack experiment with a 7B model on one A6000 GPU takes around 20 hours. Running a 16-task, 8-shot Task Haystack experiment with a 7B model on one A6000 GPU takes around 8 hours. For 34B and 70B models, we use four A6000 GPUs. For evaluations with OpenAI models, we use the Batched API. [^8] https://platform.openai.com/docs/guides/batch All experiments using OpenAI models (Table 2 bottom rows; Fig 22-22) incur a total cost of about \$8,000 at the time of writing.
Appendix B Additional Results
B.1 Scale-Task Experiments
In Table 7, we report the results in the Scale-Task setting, where we fix the number of shots per class $n_{shot}$ to be 2, and experiment with $n_{task}\in\{8,16,24,32,40,48,56,64\}$. [^9] Since each column in the table uses a different set of tasks, accuracies and pass rates from different columns are not directly comparable. We noticed that the overall pass rates are higher than those in the Scale-Shot setting (Table 2), potentially due to a smaller value of $n_{shot}$. However, long-context models still struggle in this setting: in 41 out of 44 cases in Table 7, the overall pass rates drop below 90%. The three cases achieving pass rates above 90% use the Llama2-7B (80k) model, an outlier model discussed in §5.1.
Table 7: Main Results: Fixing 2 Shots, Scaling the Number of Tasks (§B.1). See the caption of Table 2 for the explanations of the table headers.
Model 8 tasks (4k) 16 tasks (8k) 32 tasks (15k) 64 tasks (25k)
S-acc L-acc pass S-acc L-acc pass S-acc L-acc pass S-acc L-acc pass
FILM-7B (32k) 79.1 77.1 87.5 79.1 75.1 77.5 73.3 72.0 88.1 70.6 69.7 75.3
Llama-2-7B (32k) 70.1 60.7 65.0 72.8 64.5 53.8 70.6 64.5 59.4 67.1 61.2 63.1
Llama-2-7B (80k) 49.9 58.5 97.5 49.8 60.2 100.0 49.5 58.3 91.2 48.6 52.0 89.7
Llama-3-8B (1048k) 68.3 65.4 75.0 70.0 69.1 76.2 67.4 65.1 75.6 66.4 65.7 81.2
Llama-3-70B (1048k) 77.1 73.8 45.0 79.0 74.4 50.0 76.0 61.6 59.4 74.1 70.4 70.3
Llama-3.1-70B (128k) 81.5 78.2 77.5 82.8 81.1 76.2 79.2 78.0 72.5 76.9 75.6 75.3
Yi-6B (200k) 72.0 54.4 50.0 73.0 58.6 51.2 68.4 59.2 63.7 63.7 55.7 65.6
Yi-9B (200k) 78.6 73.4 62.5 77.7 72.9 71.2 75.5 70.3 61.3 70.2 66.8 61.3
Yi-34B (200k) 66.1 70.7 87.5 74.1 72.4 60.0 74.0 69.7 63.1 71.5 68.2 59.4
B.2 Task Subset with Permissive Licenses
To ensure responsible data use, in this section we introduce a new subset of 16 tasks (Table 9), each with a permissive license. We recommend that users of Task Haystack refer to this subset for future benchmarking and analysis. We have conducted evaluations on selected models with this subset. We also evaluated recent models that were released after the submission date, including meta-llama/Llama-3.2-1B-Instruct, meta-llama/Llama-3.2-3B-Instruct and gpt-4o-mini-2024-07-18. The results are presented in Table 9.
Table 8: Subset of 16 Tasks with Permissive Licenses (§B.2).
metaphor-boolean fever function-of-decision-section climate-commitments-actions
amazon-massive-scenario silicone-dyda-da brag-action student-question-categories
acl-arc wic semeval-absa-laptop senteval-cr
dbpedia14 wiki-hades environmental-claims babi-nli
Table 9: Additional Results on 16 Tasks with Permissive Licenses (§B.2). Fixing 16 Tasks, Scaling the Number of Shots. See caption of Table 2 for the explanations of the table headers. "-" indicates that the prompt exceeds the maximum context length.
Model 1-shot (4k) 2-shot (8k) 4-shot (16k) 8-shot (32k)
S-acc L-acc pass S-acc L-acc pass S-acc L-acc pass S-acc L-acc pass
Mistral-7B (32k) 67.5 68.9 95.0 70.7 69.4 70.0 72.3 69.7 52.5 - - -
FILM-7B (32k) 68.9 71.1 91.2 71.4 71.9 87.5 72.7 72.4 73.8 - - -
Llama-3.1-70B (128k) 73.8 74.2 83.8 75.0 75.1 88.7 76.6 75.7 77.5 77.5 75.6 73.8
Llama-3.2-1B (128k) 50.6 43.9 71.3 52.2 44.6 68.8 57.9 46.3 62.5 58.4 45.7 68.7
Llama-3.2-3B (128k) 59.2 62.7 93.8 60.8 63.3 62.5 64.0 64.2 68.8 63.7 64.4 66.2
Gemini-1.5-Flash (128k) 74.6 76.5 100.0 73.1 76.8 97.5 70.8 77.2 91.2 68.3 77.2 87.5
GPT-4o-mini (128k) 73.0 72.2 80.0 74.0 72.8 82.5 73.0 73.7 80.0 72.3 73.8 85.0
B.3 Controlled Analysis
In Fig. 8-8, we repeat the controlled experiments in Fig. 5-5, using N-task=64, N-shot=2 instead of N-task=16, N-shot=4. Our observations are generally consistent with those in Fig. 5-5. One exception is Mistral-7B (32k) experiments in Fig. 8, where the model achieves comparable accuracies in Recall, Replay and Remove. We attribute this to the usage of a smaller N-shot value compared to Fig. 5.
B.4 Comparing Base and Instruct Models
Previously in Table 2, we experimented with the base version of Llama-3.1-70B model (i.e., meta-llama/Llama-3.1-70B), and identified it as the most capable open-weight model among those evaluated. Here in Table 10, we additionally report results using its instruct version (i.e., meta-llama/Llama-3.1-70B-Instruct). Compared to the base model, the instruct model demonstrates improvements in S-acc across all settings, with L-acc remaining comparable or slightly improved. The increased S-acc raises the reference threshold, resulting in lower overall pass rates.
One possible explanation is that instruction tuning and subsequent post-training processes, which primarily involve shorter texts and conversational data, may encourage the model to focus more on the beginning of the context. This shift could potentially compromise its ability to handle long contexts. Given this observation, we believe Lifelong ICL and Task Haystack can also serve as a tool to monitor the long-context modeling capabilities before and after post-training processes.
Table 10: Comparing Base and Instruct Versions of Llama-3.1-70B. We use the 16-task Scale-Shot setting, consistent with Table 2.
Model 0-shot 1-shot (4k) 2-shot (8k) 4-shot (16k) 8-shot (32k)
Llama-3.1-70B (128k) 58.8 81.7 81.2 80.0 82.8 81.1 76.2 84.6 82.4 83.8 85.2 83.3 80.0
Llama-3.1-70B-Inst. (128k) 78.4 84.0 82.2 72.5 84.9 82.4 63.7 85.8 82.7 72.5 86.9 83.2 67.5
B.5 Original NIAH Experiments